a dose of methylnaltrexone bromide is administered to patients weighing over 251 pounds at a level of 0.15 mg per kilogram of body mass. each dose is administered through a subcutaneous injection every other day. if a patient weighs 264 pounds and the methylnaltrexone bromide solution has a concentration of 20 mg/ml, how many milliliters will be administered in a period of two weeks?
Answer #1

Administerd dose of methylnaltrexone bromide is = 0.15 mg/kg of body mass every other day.

Patient's weight = 264 pounds= 264/2.2= 120 kg

Hence require dose as per atients weight= 120X 0.15 mg=18 mg

Every other day signifies  day1, day3, day5, day7, etc.

Means 4 days in one week.

Hence for two weeks 8 doses will be required.= 18X 8= 144 mg

Available methylnaltrexone bromide solution concentration = 20 mg/ml

hence for 144 mg required ml= 144/20= 7.2 ml in a period of 2 weeks